two particles of mass 2kg and 1kg are moving along the same line and same direction with speeds 2 meter per second and 5 meter per second respectively. what is the speed of centre of mass of the system?

Grade:

1 Answers

Arun
6 years ago
Dear Shubhangi
 
Suppose Vc is the velocity of centre of mass of the system so we have,
 
(m1 + m2) Vc = m1 v1 + m2 v2
Vc = 9/3 = 3 m/sec.
 
Hence the velocity of centre of mass is 3 metre per second.
 
Regards
Arun (askIITians forum expert)